A system undergoes a process consisting of the following two steps

Step 1:the system absorbs 73 J of heat while 35 J of work done is done on it.

Step 2: the system absorbs 35 of j heat while performing 72 J of work

Calculate the change in internal energy for the overall process?


1
Expert's answer
2022-01-31T17:14:01-0500

Step 1


"\Delta\>u=q+w"


"=73+35=108"


Step 2


"\Delta\>u=35-72=-37"


Overall change "=108-37"

"=71J"